OpenStudy (anonymous):

What Is The 13th Term In The Fibonacci Sequence? Term 1 2 3 4 5 6 7 8 9 Fibonacci Number 1 1 2 3 5 8 13 21 34

OpenStudy (anonymous):

Next term can be found by sum of previous two terms..

OpenStudy (anonymous):

34 + 21 = 55 is the next term.. Like wise 55 + 34 = 89 is its next term.. Like wise find the 13th term here..

OpenStudy (cwrw238):

each term is formed by adding the previous 2 terms so the 10th term is 21 + 34 = 55 repeat this procedure 3 more times and you have the 13th term

OpenStudy (anonymous):

Wait I add the Fibonacci numbers? or te terms?

OpenStudy (anonymous):

Ohh Okay Thanks.

OpenStudy (anonymous):

Fibonacci..

OpenStudy (cwrw238):

get it?

OpenStudy (anonymous):

Okay I Got 142.

OpenStudy (anonymous):

So, what is the 13th term you got?

OpenStudy (anonymous):

No.. How??

OpenStudy (anonymous):

What you added??

OpenStudy (anonymous):

I added 34+55=89 Then 55+89. Thats wrong?

OpenStudy (anonymous):

55 + 89 = ??

OpenStudy (anonymous):

142

OpenStudy (anonymous):

Ohh crap Okay I See.

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

Add it carefully..

OpenStudy (anonymous):

144

OpenStudy (cwrw238):

5 + 89 = 144

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

Yep.. Do it one more time..

OpenStudy (anonymous):

*55??

OpenStudy (anonymous):

233

OpenStudy (anonymous):

You got this as your 12th term, do one more time, you will get 13th term.. Yep you are right.. :)

OpenStudy (anonymous):

Thank you!

OpenStudy (cwrw238):

sorry i meant 55 + 89

OpenStudy (anonymous):

It is okay @cwrw238

OpenStudy (anonymous):

You are welcome dear..

OpenStudy (anonymous):

Okay So The Next Question Asks What is the sum of the first 13 terms in the Fibonacci Sequence? What Does That Mean?

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@BeastMuch it's 233 btw

OpenStudy (anonymous):

tht 13 number

OpenStudy (anonymous):

Ohh So Its The Same Ting??

OpenStudy (anonymous):

yep :]

OpenStudy (anonymous):

Okay So What About This One? What is the ratio between the 13th and 12th terms in the Fibonacci Sequence? Round your answer to five decimal places.

OpenStudy (anonymous):

umm not to sure bout tht one you'll have to give me a minute :)

OpenStudy (anonymous):

OKay Thanks.

OpenStudy (cwrw238):

12th term is 144 and 13th term is 233 the ratio is 233 / 144 which you can find on your calculator

OpenStudy (anonymous):

Would You Divide?

OpenStudy (cwrw238):

the sum of the first 13 terms is what it says 1+ 1 + 2 + ..... +144+ 233

OpenStudy (cwrw238):

233 divided by 144

OpenStudy (anonymous):

1.61805 ?

OpenStudy (anonymous):

yaw

OpenStudy (cwrw238):

1.61806 is more exact but they would probably accept 233/144 as the answer - just write them both !!!

OpenStudy (anonymous):

Haha OKay Thanks.

OpenStudy (cwrw238):

yw